OpenStudy (anonymous):

sasha is arranging ride so that 27 of her members can go blowing ,some people can ride and vans and the rest must ride in cars . if 12 go in vans and 5 rides in each car how many cars will they need?

OpenStudy (amistre64):

ive never tried blowing a bowling ball down the alley

OpenStudy (s):

12 + 5x = 27 5x = 27-12 5x = 15 x = 3 cars